OpenStudy (anonymous):

Solve: | x/3 +1 | = 6 {-7, 15} {-21, 15} {7, 15} {-15, 21} To clarify, the x/3 is a fraction, not a division sign

OpenStudy (anonymous):

y do you like absolute value so much ( i hate it xD )

OpenStudy (anonymous):

I don't lol homework.......

OpenStudy (anonymous):

it happens ^^

OpenStudy (anonymous):

lol sorries i read the problem wrong

OpenStudy (anonymous):

lemme try it again :P

OpenStudy (anonymous):

can you answer this question?

OpenStudy (anonymous):

ya should be able to

OpenStudy (anonymous):

i just thought the 1 was part of the denom.

OpenStudy (anonymous):

{-21, 15} I actually got it thanks for the effort though!:)

OpenStudy (anonymous):

|x/3+1|=6 (-1)+x/3+1=6 +(-1) additive inverse x/3 = 6 -1 simplify 3 [ x/3=5 ] multiply by reciprocal x = 15 or |x/3 +1| = 6 (-21)/ 3 + 1 =6 |-7 +1|=6 |-6|+6 6= 6